# The fundamental localization phases in quasiperiodic systems: A unified framework and exact results

### v2

@@ III.2 Universal mechanism for the emergence of critical states
We now show a new universal mechanism for the critical states in spinful QP chains: Spinful quasiperiodic systems can host critical states if the system possesses generalized incommensurate matrix element zeros ${\cal G}_{\Pi}$, which are defined as
${\cal G}_{\Pi}=\Big{\{}j_{k}\Big{|}\lim_{L\rightarrow\infty}p_{j_{k}}^{s}=0,\;s\in\{0,x,y,z\}\Big{\}}.$ (9)
This is a novel generalization of the mechanism for critical states in spinless QP chains to the spinful systems. The generalized incommensurate zeros (GIZs) in matrix elements refer to the incommensurately distributed zeros (IDZs) located in any component of the hopping coupling matrix $\Pi_{j}$ in the thermodynamic limit. The IDZs is defined as the site indices $\{j_{1},j_{2},\dots\}$ at which the hopping coefficients vanish in the thermodynamic limit ${\cal G}=\{j_{k}\mid\lim_{L\rightarrow\infty}t_{j_{k}}=0\}$ (Zhou et al., 2023; Simon and Spencer, 1989; Jitomirskaya and Marx, 2012), and the distribution of $\{j_{k}\}$ is incommensurate with the periodic of the system size. Similarly, the presence of GIZs in spinful models partitions the system into segments that remain invariant under dual transformations. This dual-invariance structure drives delocalized states into critical states [Fig. 2(b)].
The proof of this mechanism relies on the invariance of critical states under dual transformations. In the presence of the GIZs, the system includes processes $V_{j}^{\mathrm{od}}c_{j+1s}^{\dagger}c_{js^{\prime}}+\mathrm{h.c.}$, where $s$ and $s^{\prime}$ can be either identical or different. Under the dual transformation, these terms remain invariant as $V_{n}^{\mathrm{od}}c_{n+1s}^{\dagger}c_{ns^{\prime}}+\mathrm{h.c.}$ Consequently, when these terms dominate, critical orbitals enter the spectrum. From the perspective of the parent Hofstadter Hamiltonian, the spin-1/2 QP chain is mapped to a charged particle moving in a 2D bilayer system under a magnetic field. In this mapping, QP hopping in the hopping coupling matrix corresponds to diagonal hopping within or between the layers. This bilayer diagonal motion contributes dispersions along both $x$- and $y$-directions, thus introducing critical states into the projected 1D spectrum.
Inspired by this proof based on invariance under the dual transformation, we propose an additional guiding principle for generating critical states in spinful QP chains: introducing IDZs in the component of the on-site matrix shared with the hopping coupling matrix. For example, if the hopping coupling matrix are uniform, $\Pi_{j}\sim t\sigma_{+}$, then the system exhibits critical states if the on-site matrix $M_{j}$ exhibit IDZs in either the $\sigma_{x}$ or $\sigma_{y}$ component. This principle is demonstrated by the emergence and absence of critical states in mosaic lattice models and the new exactly solvable models we introduce in the next section. Notably, this principle is unique to spinful QP chains and does not extend to spinless QP chains. The key difference is that multiple generators in spinful QP chain allow for the transfer of IDZs from $M_{j}$ into either energy-dependent unbounded on-site potentials or IDZs in the hopping matrix, thereby generating critical states when reducing the spinful QP chain to a single chain. Some examples are provided in Theorem III in the next subsection.
The generalized incommensurate zeros (GIZs) ${\cal G}_{\Pi}$, and the incommensurately distributed zeros (IDZs) in the shared component of on-site and hopping coupling matrices serve as two guiding principles for introducing critical orbitals into the spectrum. These principles, together with the criteria for pure phases, provide powerful insights for constructing spinful QP models that host seven fundamental localization phases. For a generic spinful quasiperiodic system, Avila’s theory cannot analytically characterize the system. This motivates us to explore the conditions under which the spinful QP system becomes exactly solvable, as discussed in the following subsection.
@@ III.3 Exact solvability of QP systems from local constraint
We now introduce the local constraint, with which the spinful QP system can exhibit exact solvability: A spinful quasiperiodic system can have exactly solvable points if the hopping coupling matrix $\Pi_{j}$ or its dual counterpart $\Pi_{n}$ are degenerate, which corresponds to
@@
together with $p_{j}^{s}$ and off-diagonal terms of $M_{j}$ being either constant or purely quasiperiodic. This condition represents a local constraint, where the spinful quasiperiodic systems can be reduced to an effectively 1D spinless QP chain for the dressed particle with only nearest neighbor hopping, as illustrated in Fig. 2(c). In this case, Avila’s global theory can be applied to analytically characterize the localization properties of the dressed particles.
To elaborate the physics of this universal result, we consider the generic eigen-equation for the spinful QP system [Eq. (1)] $H\lvert\Psi\rangle=E\lvert\Psi\rangle$, with $\lvert\Psi\rangle=\sum_{j=1}^{L}(c^{\dagger}_{j,\uparrow}u_{j,\uparrow}+c^{\dagger}_{j,\downarrow}u_{j,\downarrow})\lvert\mathrm{vac}\rangle$, which is given by
@@
where $\vec{u}_{j}=(u_{j\uparrow},u_{j\downarrow})^{\intercal}$ denotes the spinnor for spin-1/2 particles. Under the constraint described in Eq.(10), the hopping coupling matrix takes one of the following fundamental forms: for spin-conserved processes, $\tilde{\Pi}_{j}\sim(\sigma_{0}\pm\sigma_{z})/2=\Lambda_{\pm}$; for spin-flipped processes, $\tilde{\Pi}_{j}\sim(\sigma_{x}\pm i\sigma_{y})/2=\sigma_{\pm}$; or a combination of both processes. Such constraint reduces the otherwise complicated eigen-equation in Eq.(11) to a 1D spinless QP chain with nearest-neighbor hopping. Here, $\tilde{\Pi}_{j}$ represents the hopping coupling matrix after a local unitary transformation, and the transformed spinnor is denoted by $\vec{v}_{j}$. We present the essential physics for the first two fundamental cases in the main text, and additional details for all three cases can be found in Appendix B.
@@
with effective hopping coefficient and on-site potential given by
$t_{j}^{\mathrm{eff}}=\frac{t_{j}M_{j}^{12}}{E-M_{j}^{22}},\quad V_{j}^{\mathrm{eff}}=M_{j}^{11}+\frac{t_{j-1}^{2}}{E-M_{j-1}^{22}}+\frac{M_{j}^{12}M_{j}^{21}}{E-M_{j}^{22}}.$ (17)
These effective hopping and on-site potentials in Eqs. (14) and (17) highlight the requirement for $p_{j}^{s}$ and the off-diagonal elements of $M_{j}$ to be either constant or purely quasiperiodic in order to preserve exact solvability, as outlined in Theorem III (Sec. III.3). The effective eigen-equation for the dressed particles involves processes such as $M_{j}^{12}M_{j}^{21}$, $t_{j}M_{j}^{12}$, and $t_{j-1}^{2}$. These terms maintain a single QP frequency modulation when $p_{j}^{s}$ and off-diagonal elements of $M_{j}$ are either constant or purely quasiperiodic. However, when these terms are a mixture of constant and QP components, such as in the form $[A+B\cos(2\pi\alpha j)]^{2}$, they lead to mixed frequency modulations, thereby breaking the exact solvability.
The effective hopping coefficient and on-site potential in Eq. (17) explain the guiding principle from theorem II (Sec. III.2), where critical states arise if the system possesses IDZs in the shared components of both coupling matrices. Specifically, for the hopping coupling matrix $\Pi_{j}\sim t\sigma_{+}$, the presence of IDZs in $M_{j}^{12}$, which couples to $\sigma_{x}$ and/or $\sigma_{y}$, ensures that the effective 1D system exhibits IDZs, thus facilitating the emergence of critical states. This behavior is exemplified by the absence (or presence) of critical states in type-I (or type-II) mosaic lattice models Wang et al. (2020a); Zhou et al. (2023). For a comprehensive introduction to mosaic lattice models and their unification within the spin-1/2 QP framework, see Appendix C. The various types of mosaic lattice models can be unified as
where $c_{j}=(c_{j,\uparrow},c_{j,\downarrow})^{\intercal}$ is the spinor for the annihilation operators, $\lambda$ is the uniform hopping strengths. The onsite matrix $V_{j}^{\mathrm{M}}$ distinguishes the type-I and type-II mosaic lattices, respectively given by
@@
with $V_{0}$ and $t$ being the strength of QP modulation of the on-site matrix for type-I and type-II mosaic models, respectively. In the type-I mosaic lattice model, there are no IDZs in the shared components, resulting in a spectrum composed solely of extended and localized states Wang et al. (2020a). In contrast, for the type-II mosaic lattice model, IDZs appear on $\sigma_{x}$, which introduces energy-dependent quasiperiodic hopping and an unbounded on-site potential, thereby giving rise to rigorously defined critical states within the spectrum Zhou et al. (2023).
@@
The universal results offer a powerful guidance for constructing exactly solvable models hosting various localization physics. The effective hopping and potential terms [Eq. (14)-(17)] provide a framework for designing the microscopic details of these systems by controlling the energy and coupling between the internal degrees of freedom. For example, by designing the effective potential, one can induce or suppress critical states, as well as manipulate the localization length.
@@
Figure 3: New exactly solvable models constructed using universal results and dual transformations. The fractal dimension (FD) as a function of energy $E$ and hopping strengths $\lambda/t$. (a) The quasiperiodic spin-flipped (QPSF) model, derived by removing the $\sigma_{0}$ component of the type-II mosaic model, exhibits pure localized and critical phases. (b) The dual counterpart of the type-II mosaic lattice model, featuring analytic mobility edges (MEs) that separate extended and critical states. The MEs occur at $E_{c}=\pm\lambda$ and are marked by solid lines. (c) The dual counterpart of the QPSF model, which is equivalent to removing the $\sigma_{0}$ component in the dual model from (b). The dual-QPSF model exhibits pure extended and critical phases, with the phase transition points $\lambda=t$ marked by dashed lines. The system size is $L=2586$.
III.4 Examples of exactly solvable models based on universal results
@@
With the guidance from the universal results, we introduce several new exactly solvable models derived from the type-II mosaic lattice model (Zhou et al., 2023) in Eq. (18) and Eq. (20) by removing the MEs. Through a combination of dual transformation, we can construct new nontrivial models, which demonstrate the application of the theorems. We investigate the phase diagram of the proposed model both analytically and numerically. We numerically solve the spectrum and employ the fractal dimension (FD) to phenomenologically characterize the localization properties of the eigenstates $\lvert\Psi\rangle=\sum_{j=1}^{L}u_{j}a_{j}^{\dagger}\lvert\mathrm{vac}\rangle$, where $\mathrm{FD}=-\lim_{L\rightarrow\infty}\ln\sum_{j=1}^{L}|u_{j}|^{4}/\ln L$, with $u_{j}$ being the wave function coefficients, $L$ the system size. In 1D, the FD approaches 1 for extended states and 0 for localized states, while for critical states, $10<\mathrm{FD}<$ 1. The FD quantifies the effective dimension experienced by an eigenstate: extended states uniformly spread across the system, yielding $\mathrm{FD}=1$. Localized states decay exponentially, effectively perceiving zero dimension, hence $\mathrm{FD}=0$. Critical states, however, exhibit self-similar wave function structures that span the entire system, resulting in an effective dimension between 0 and 1.
@@
We begin with the type-II mosaic lattice model (Zhou et al., 2023) in Eq. (18) and Eq. (20), whose Hamiltonian is given by
@@
where $\lambda$ is the uniform hopping strengths and $t$ is the strength of balanced QP potential and exchange coupling. The spectrum comprises critical (localized) states for the energies satisfy $E<|\lambda|$ ($E>|\lambda|$(Zhou et al., 2023).
Next, we construct new exactly solvable models that exhibit phase transitions between pure critical and localized phases by eliminating the MEs, as outlined in theorem I (Sec. III.1). This is accomplished by removing the spin-independent component of the on-site matrix. The resulting quasiperiodic spin-flipped (QPSF) model is given by
@@
This behavior can be understood as follows: In the limit $\lambda\rightarrow 0$, the QP on-site term localizes the system into dimmers with eigenenergies $E=\pm 2tV_{j}^{\rm d}$. In this case, the energy difference between adjacent site fluctuates significantly, and turning on the unifrom hopping $\lambda$ only hybridize a few dimmers with close energies, resulting in a localized phase. Conversely, in the limit $t\rightarrow 0$, the unifrom spin-flipped hopping creates degenerate dimers with flat band energies $E=\pm\lambda$. The degeneracy of these dimers is then lifted by the QP on-site term $V_{j}^{\rm d}$, which hybridizes them drastically, leading to a delocalized phase. The delocalized states are guaranteed to be critical due to the IDZs in the effective hopping
@@
Furthermore, we develop new models by performing dual transformations to the two aforementioned models, which exhibit mobility edges or phase transitions between critical and extended states. Applying the dual transformation to the type-II mosaic model in Eq. (21) results in the following dual counterpart
@@
$\tilde{T}_{n}=2t\Lambda_{+},\quad\tilde{M}_{n}=\lambda\big{[}\cos(2\pi\alpha n)\sigma_{z}+\sin(2\pi\alpha n)\sigma_{y}\big{]}.$ (25)
In the limit $t\rightarrow 0$, the system yields two flat bands from the on-site matrix with energies $E=\pm\lambda$. The quantum states within these flat bands can be combined into either localized, extended, or critical states. The uniform hopping drives some states into extended states, while the IDZs in the shared component lead other states into critical states.
Similarly, the dual model of the QPSF model in Eq. (22) is given by
@@
We have demonstrated a direct application of the universal results to both explain existing models and develop new nontrivial ones. In the following section, we apply these theorems to construct novel models that incorporate all fundamental localization phases and MEs, providing a comprehensive phase diagram for the QP optical Raman lattice model.

### v4

@@ III.2 Universal mechanism for the emergence of critical states
We further show a new universal mechanism for the critical states in spinful QP chains: Spinful quasiperiodic systems can host critical states if the system possesses generalized incommensurate matrix element zeros ${\cal G}_{\Pi}$, which are defined by
${\cal G}_{\Pi}=\Big{\{}j_{k}\Big{|}\lim_{L\rightarrow\infty}\Pi_{j_{k}}^{\alpha\beta}=0\Big{\}},$ (9)
where $\alpha,\beta\in\{1,2\}$ label the matrix element indices. The generalized incommensurate zeros (GIZs) in the above equation refer to the IDZs for any component of the hopping coupling matrix $\Pi_{j}$ in the thermodynamic limit. This is a nontrivial generalization of the mechanism for critical states in spinless QP chains Zhou et al. (2023) to the present spinful systems. In spinless systems, the IDZs partition the system into multiple segments, driving the delocalized wave function to reorganize and develop the characteristic self-similar structure of critical states Zhou et al. (2023); Liu (2024). In the current spinful systems, the dual-invariant GIZs will similarly partition system into incommensurately distributed segments in the internal degree of freedom (spin) and the real space, leading to critical states with multifractal features [Fig. 2(b)]. When the QP hopping matrix $\Pi_{j}$ with GIZs dominates over the on-site QP couplings $M_{j}$, the critical states emerge.
We prove this result by mapping the spinful QP model to a generalized 2D bilayer square lattice system, in the presence of an external magnetic field with an irrational flux threading each unit cell. Under this mapping, the total Hamiltonian Eq. (1) becomes (see Appendix C)
$H=$ $\sum_{x,y,s,s^{\prime}}\Big{[}t^{ss^{\prime}}a_{x+1,y,s}^{\dagger}a_{x,y,s^{\prime}}+\frac{V^{ss^{\prime}}}{2}e^{-i\varphi x}a_{x,y+1,s}^{\dagger}a_{x,y,s^{\prime}}$
$+\lambda^{ss^{\prime}}a_{x,y,s}^{\dagger}a_{x,y,s^{\prime}}+\frac{\mu^{ss^{\prime}}}{2}(e^{-i\varphi x}a_{x+1,y+1,s}^{\dagger}a_{x,y,s^{\prime}}$
$+e^{i\varphi x}a_{x+1,y-1,s}^{\dagger}a_{x,y,s^{\prime}})+{\rm h.c.}\Big{]},$ (10)
where $\varphi=2\pi\alpha$, $s,s^{\prime}$ denote the layer indices, $t^{ss^{\prime}}$ and $\mu^{ss^{\prime}}$ are the uniform and QP hopping coefficients of elements in $\Pi_{j}$, respectively, and $\lambda^{ss^{\prime}}$ and $V^{ss^{\prime}}$ are the uniform and QP on-site coefficients of elements in $M_{j}$, respectively. The QP hopping term $\mu^{ss^{\prime}}V_{j}^{\mathrm{od}}c_{j+1s}^{\dagger}c_{js^{\prime}}+\mathrm{h.c.}$ with GIZs in the Hamiltonian Eq. (1) characterizes the intra-layer (for $s=s^{\prime}$) or inter-layer (for $s\neq s^{\prime}$) hopping coupling in the diagonal direction $(x+1,y\pm 1)\leftrightarrow(x,y)$. In contrast, the uniform hopping matrix elements and onsite terms are mapped to the corresponding couplings along $x$ and $y$ directions, respectively. When the QP hopping dominates over other terms, e.g. when $\mu^{ss^{\prime}}\gg t^{ss^{\prime}},\lambda^{ss^{\prime}},V^{ss^{\prime}}$, the cyclotron motion must be extended in the diagonal direction, leading to delocalization in both $x$ and $y$ directions (the real space and dual space). The dual transformation is equivalent to a gauge transformation in 2D bilayer (see Appendix C for details)
$a^{\dagger}_{x,y,s}\rightarrow e^{i\varphi xy}a^{\dagger}_{x,y,s},\ \vec{A}=(0,-\varphi x)\rightarrow\vec{A}=(\varphi y,0),$ (11)
which clearly preserves the delocalization of cyclotron orbits in the two directions, rendering the invariance under the dual transformation. Thus these states exhibit self-duality, as the key signature of critical states.
In addition to the GIZs, we present a secondary theorem for rigorous realizing critical states in spinful QP chains by connecting the spinful system to the spinless QP models. Specifically, when the on-site matrix $M_{j}$ has IDZs in the matrix component which is shared by the hopping term $\Pi_{j}$, one can reduce the spinful QP chain to an effective 1D spinless model, with the IDZs in $M_{j}$ being transferred into either IDZs in the hopping coefficients or into incommensurately distributed divergence points in the on-site potentials. We therefore reduce to the the universal mechanism in the spinless system for generating the critical states. For instance, when $\Pi_{j}\sim t\sigma_{+}$ is uniform, critical states emerge if $M_{j}$ host IDZs in its $\sigma_{x}$ or $\sigma_{y}$ coefficients. We shall show this principle through the quasiperiodic mosaic models and the new exactly solvable models proposed in the next subsection.
The GIZs ${\cal G}_{\Pi}$ and the IDZs in the matrix component of $M_{j}$ shared in $\Pi_{j}$ serve as two powerful guiding principles to rigorously realize the critical states. These two principles can also be proved in an intuitive way based on the semiclassical approximation (see Appendix D), which illuminates the origin of the critical states. Together with the criteria for pure phases, these principles provide insights into construction of spinful QP models that host all the fundamental localization phases.
@@ III.3 Exact solvability of QP systems from local constraint
We now present the third universal result regarding exact solvability of the spinful QP system: A spinful quasiperiodic system can have exactly solvable points if the hopping coupling matrix $\Pi_{j}$ or its dual counterpart $\Pi_{n}$ are degenerate, which corresponds to
@@
together with $p_{j}^{s}$ and off-diagonal terms of $M_{j}$ being either constant or purely quasiperiodic. This condition represents a local constraint, where the spinful QP systems can be reduced to an effectively 1D spinless QP chain for the dressed particle with only nearest-neighbor hopping, as illustrated in Fig. 2(c). In this case, Avila’s global theory can be applied to analytically characterize the localization properties of the dressed particles.
To elaborate the physics of this universal result, we consider the generic eigen-equation for the current spinful QP system that $H\lvert\Psi\rangle=E\lvert\Psi\rangle$, with $\lvert\Psi\rangle=\sum_{j=1}^{L}(c^{\dagger}_{j,\uparrow}u_{j,\uparrow}+c^{\dagger}_{j,\downarrow}u_{j,\downarrow})\lvert\mathrm{vac}\rangle$, which is given by
@@
where $\vec{u}_{j}=(u_{j\uparrow},u_{j\downarrow})^{\intercal}$ denotes the spinor for spin-1/2 particles. Under the constraint described in Eq. (12), the hopping coupling matrix takes one of the following fundamental forms: for spin-conserved processes, $\tilde{\Pi}_{j}\sim(\sigma_{0}\pm\sigma_{z})/2=\Lambda_{\pm}$; for spin-flipped processes, $\tilde{\Pi}_{j}\sim(\sigma_{x}\pm i\sigma_{y})/2=\sigma_{\pm}$; or a combination of both processes. Such constraint reduces the otherwise complicated eigen-equation in Eq. (13) to an effective 1D spinless QP chain with nearest-neighbor hopping. Here, $\tilde{\Pi}_{j}$ represents the hopping coupling matrix after a local unitary transformation, and the transformed spinor is denoted by $\vec{v}_{j}$. We present the essential physics for the first two fundamental cases in the main text, and additional details for all three cases can be found in Appendix E.
@@
with effective hopping coefficient and on-site potential
$t_{j}^{\mathrm{eff}}=\frac{t_{j}M_{j}^{12}}{E-M_{j}^{22}},\ V_{j}^{\mathrm{eff}}=M_{j}^{11}+\frac{t_{j-1}^{2}}{E-M_{j-1}^{22}}+\frac{M_{j}^{12}M_{j}^{21}}{E-M_{j}^{22}}.$ (19)
The effective hopping and on-site potentials in Eqs. (16) and (19) highlight the requirement for $p_{j}^{s}$ and the off-diagonal elements of $M_{j}$ to be either constant or purely quasiperiodic in order to preserve exact solvability, as outlined in Theorem III (Sec. III.3). The effective eigen-equation for the dressed particles involves processes such as $M_{j}^{12}M_{j}^{21}$, $t_{j}M_{j}^{12}$, and $t_{j-1}^{2}$. These terms maintain a single QP frequency modulation when $p_{j}^{s}$ and off-diagonal elements of $M_{j}$ are either constant or purely quasiperiodic. However, when these terms are a mixture of constant and QP components, such as in the form $[A+B\cos(2\pi\alpha j)]^{2}$, they lead to mixed frequency modulations, thereby breaking the exact solvability.
The effective hopping coefficient and on-site potential in Eq. (19) also explain the guiding principle from theorem II (Sec. III.2), where critical states arise if the system possesses IDZs in the shared components of both coupling matrices. Specifically, for the hopping coupling matrix $\Pi_{j}\sim t\sigma_{+}$, the presence of IDZs in $M_{j}^{12}$, which couples to $\sigma_{x}$ and (or) $\sigma_{y}$, ensures that the effective 1D system exhibits IDZs, hence leading to the emergence of critical states. This result is exemplified by the absence (or presence) of critical states in type-I (or type-II) QP mosaic models Wang et al. (2020a); Zhou et al. (2023). For a unified characterization, we describe the both types of quasiperiodic mosaic models in the spin-$1/2$ QP framework, which can be written as (see more details in Appendix F)
$H_{\mathrm{M}}=\sum_{j}\lambda(c^{\dagger}_{j+1}\sigma_{-}c_{j}+\mathrm{h.c.})+\sum_{j}c_{j}^{\dagger}V_{j}^{\mathrm{M}}c_{j}.$ (20)
Here $c_{j}=(c_{j,\uparrow},c_{j,\downarrow})^{\intercal}$ is the spinor for the annihilation operators, and $\lambda$ is the uniform hopping strength. The onsite matrix $V_{j}^{\mathrm{M}}$ distinguishes the type-I and type-II mosaic lattices, respectively given by
@@
with $V_{0}$ and $t$ being the strength of QP modulation of the on-site matrix for type-I and type-II mosaic models, respectively. In the type-I quasiperiodic mosaic model, there are no IDZs in the shared components, resulting in a spectrum composed solely of extended and localized states Wang et al. (2020a). In contrast, for the type-II quasiperiodic mosaic model, IDZs appear on $\sigma_{x}$, which introduces energy-dependent quasiperiodic hopping and an unbounded on-site potential, thereby giving rise to rigorously defined critical states within the spectrum Zhou et al. (2023).
@@
The universal results offer a powerful guidance to construct exactly solvable models hosting all types of localization physics. The effective hopping and onsite coupling terms [Eq. (16)-(19)] provide a framework for designing the microscopic details of these systems by controlling the energy and coupling between the internal degrees of freedom. For example, by designing the effective potential, one can induce or suppress critical states, as well as manipulate the localization length.
@@
Figure 3: New exactly solvable models constructed using universal results and dual transformations. The fractal dimension (FD) versus energy $E$ and hopping strengths $\lambda/t$. (a) The quasiperiodic spin-flipped (QPSF) model, obtained by removing the $\sigma_{0}$ component of the type-II QP mosaic model, exhibits pure localized and critical phases. (b) The dual counterpart of the type-II QP mosaic model, featuring analytic mobility edges (MEs) that separate extended and critical states. The MEs occur at $E_{c}=\pm\lambda$ and are marked by solid lines. (c) The dual QPSF model, obtained by removing the spin-independent component in the dual model from (b). It exhibits pure extended and critical phases, with the phase transition points $\lambda=t$ marked by dashed lines. The system size is $L=2586$.
III.4 New exactly solvable mosaic models
@@
With the universal results presented above, we put forward several new exactly solvable models derived from the type-II QP mosaic model (Zhou et al., 2023) in Eq. (20) and Eq. (22) by removing the MEs. Through a combination of dual transformation, we can construct new nontrivial models, which demonstrate the applications of the above results, and investigate the phase diagram of the models both analytically and numerically. We numerically solve the spectrum and employ the fractal dimension (FD) to phenomenologically characterize the localization properties of the eigenstates $\lvert\Psi\rangle=\sum_{j=1}^{L}u_{j}a_{j}^{\dagger}\lvert\mathrm{vac}\rangle$, where $\mathrm{FD}=-\lim_{L\rightarrow\infty}\ln\sum_{j=1}^{L}|u_{j}|^{4}/\ln L$, with $u_{j}$ being the wave function coefficients, $L$ the system size. In 1D, the FD approaches $1$ for extended states and 00 for localized states, while for critical states, $10<\mathrm{FD}<$ 1. The FD quantifies the effective dimension experienced by an eigenstate: extended states uniformly spread across the system, yielding $\mathrm{FD}=1$. Localized states decay exponentially, effectively perceiving zero dimension, hence $\mathrm{FD}=0$. Critical states, however, exhibit self-similar wave function structures that span the entire system, resulting in an effective dimension between 00 and $1$.
@@
We begin with the type-II QP mosaic model (Zhou et al., 2023) in Eq. (20) and Eq. (22), whose Hamiltonian is given by
@@
where $\lambda$ is the uniform hopping strengths and $t$ is the strength of balanced QP potential and exchange coupling. The spectrum comprises critical (localized) states for the energies satisfy $E<|\lambda|$ ($E>|\lambda|$(Zhou et al., 2023). Next, we construct new exactly solvable models that host pure critical and localized phases by eliminating the MEs, as outlined in theorem I (Sec. III.1). This is accomplished by removing the spin-independent component of the on-site matrix of the type-II QP mosaic model. This results in a quasiperiodic spin-flipped (QPSF) model given by
@@
The physics can be understood as follows: In the limit $\lambda\rightarrow 0$, the QP on-site term localizes the system into dimmers with eigenenergies $E=\pm 2tV_{j}^{\rm d}$. In this case, the energy difference between adjacent site fluctuates significantly, and turning on the uniform hopping $\lambda$ only hybridize a few dimmers with close energies, giving a localized phase. Conversely, in the limit $t\rightarrow 0$, the uniform spin-flipped hopping creates degenerate dimers with flat band energies $E=\pm\lambda$. The degeneracy of these dimers is then lifted by the QP on-site term $V_{j}^{\rm d}$, which hybridizes them drastically, leading to a delocalized phase. The delocalized states are guaranteed to be critical due to the IDZs in the effective hopping
@@
Furthermore, we can develop new models by performing dual transformations to the two aforementioned models, which exhibit mobility edges or phase transitions between critical and extended states. Applying the dual transformation to the type-II QP mosaic model in Eq. (23) results in the following dual counterpart
@@
$\tilde{T}_{n}=2t\Lambda_{+},\ \tilde{M}_{n}=\lambda\big{[}\cos(2\pi\alpha n)\sigma_{z}+\sin(2\pi\alpha n)\sigma_{y}\big{]}.$ (27)
In the limit $t\rightarrow 0$, the system yields two flat bands from the on-site matrix with energies $E=\pm\lambda$. The quantum states of the flat bands can be combined into either localized, extended, or critical states. The extended states and critical states are driven by the uniform hopping and the IDZs in the shared component, respectively.
Similarly, the dual QPSF model can be obtained by